Suppose cells have an electromotive force of 14 V.if a resistor of 20 is connected across the cells,the terminal potential difference is 12 V.Determine the current through the 20 resistor and the internal resistor of the cells.


1
Expert's answer
2021-09-30T10:14:01-0400

Internal Resistance = (EMF-V)/L

Internal Resistance =(14-12)/20

Internal Resistance = 2/20

Internal Resistance= 0.1


Current flowing through the resistor = (EMF/Resistance)

Current flowing through the resistor = 14/0.1

Current flowing through the resistor = 140